Hello!
I need help for a geometry question. To figure out the leg of a right triangle when hypotenuse equals 10, and one given leg is 8. Thanks!
Zeshawn

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

There are two ways.
You should recognize it as multiple of the 3,4,5 right triangle.
10^2=8^2+b^2
b^2=36
b=6

Answer by stanbon(75887)

figure out the leg of a right triangle when hypotenuse equals 10,
and one given leg is 8.
------
Let the unknown leg length be "x"::
x = sqrt[10^2-8^2]
=====
x = sqrt[100-64] = sqrt(36) = 6
-------------------
Cheers,
Stan H.